# Towards Constraint-aware Learning for Resource Allocation in NFV-enabled Networks

### v1

@@ Abstract
Virtual Network Embedding (VNE) is a challenging combinatorial optimization problem that refers to resource allocation associated with hard and multifaceted constraints in network function virtualization (NFV). Existing works for VNE struggle to handle such complex constraints, leading to compromised system performance and stability. In this paper, we propose a CONstraint-Aware Learning framework for VNE, named CONAL, to achieve efficient constraint management. Concretely, we formulate the VNE problem as a constrained Markov decision process with violation tolerance. This modeling approach aims to improve both resource utilization and solution feasibility by precisely evaluating solution quality and the degree of constraint violation. We also propose a reachability-guided optimization with an adaptive reachability budget method that dynamically assigns budget values. This method achieves persistent zero violation to guarantee the feasibility of VNE solutions and more stable policy optimization by handling instances without any feasible solution. Furthermore, we propose a constraint-aware graph representation method to efficiently learn cross-graph relations and constrained path connectivity in VNE. Finally, extensive experimental results demonstrate the superiority of our proposed method over state-of-the-art baselines. Our code is available at https://github.com/GeminiLight/conal-vne.
@@ 1 Introduction
Network Function Virtualization (NFV) is a promising technique that facilitates the deployment of multiple Virtual Networks (VNs) tailored to user network demands within a shared Physical Network (PN) infrastructure (Yi et al., 2018). It is vital for domains such as cloud computing and edge computing, where dynamic and efficient resource management is essential (Zhuang et al., 2020). Virtual Network Embedding (VNE), a fundamental resource allocation problem in NFV, is critical for maintaining high Quality of Service (QoS). This process, involving the mapping of VNs to PNs, represents a significant challenge. It is an NP-hard Combinatorial Optimization Problem (COP) characterized by intricate and hard constraints (Rost & Schmid, 2020).
@@
Traditional solutions to VNE range from exact to heuristic methods often struggle with either excessive computation times or limited performance in complex network scenarios (Zhang et al., 2018; Dehury & Sahoo, 2019; Fan et al., 2023). Recently, Reinforcement Learning (RL) has been a potential direction for VNE, which learns effective solving policies with unlabeled datasets. Typically, existing RL-based methods solve the VNE problem as a Markov Decision Process (MDP)  (Haeri & Trajković, 2017; Yan et al., 2020; Zhang et al., 2023b). They build feature extractors with various neural networks, and then learn policies that iteratively select a physical node to place each virtual node until the solution is completed or constraints are violated. However, in the RL framework, due to the hard constraints of VNE, a persistent zero-violation is required at each decision timestep. This strict adherence to constraints often results in numerous failure samples where constraints are violated. For failure samples, existing studies (e.g., Yao et al. (2020); Zhang et al. (2022)) consider them as noisy and only train with data that violate no constraint; or others, like Yan et al. (2020) and He et al. (2023a), consider fixed penalties to them and discourage violations.
@@
Although these RL-based algorithms (Zhang et al., 2022; He et al., 2023a; Zhang et al., 2023b) have shown efficacy, they still suffer from several significant problems in handling complex constraints of VNE. Firstly, ignoring failure samples makes policies prone to violating critical constraints, while employing fixed penalties in reward signals does not accurately reflect the severity of constraint violations. Thus, these methods underestimate valuable sample information and hamper the learning of constraint-aware policies, which results in low feasibility guarantees. Secondly, more seriously, it is hard to avoid to encounter unsolvable instances whose feasible sets are empty in practical scenarios, due to insufficient physical resource availability or excessive virtual resource requests. It is impractical to distinguish solvable and unsolvable instances, since checking the instance solvability of an NP-hard problem is time-consuming. Failure samples caused by unsolvable instances further complicate the constraint learning process, and negatively impact the stability of training and policy performance. We have conducted a preliminary study to highlight the negative impact of unsolvable instances on training, please see Appendix C. Thirdly, VNE constraints are complex and multifaceted, involving cross-graph status interactions and bandwidth-constrained path connectivity assessments, which are not adequately captured by the feature extractors used in current studies. For a more comprehensive discussion on related work, please refer to Appendix A.
@@
To address these challenges, we propose a CONstraint-Aware Learning framework for VNE, named CONAL, achieving high solution feasibility guarantee and training stability. Concretely, to optimize performance while enhancing constraint satisfaction, we formulate the VNE problem as Constrained MDP (CMDP) (Altman, 2021). However, in the process of solution construction, if any constraints are violated, the process will be early terminated and lead to incomplete solutions. This challenges precise measurement of both the quality of solutions and the degree of constraint violations. Thus, we develop a violation-tolerant mapping method to ensure complete solution construction and a measurement function to precisely evaluate constraint violations. Additionally, to achieve persistent zero constraint violation required by VNE, we introduce a reachability analysis into the optimization objective. During training, due to the existence of unsolvable instances whose constraints are impossibly satisfied, we propose an adaptive reachability budget method to make the policy optimization more stable. It dynamically decides the violations caused by a surrogate policy as budgets based on the instance’s solvability, rather than always setting budgets to zero. Furthermore, to finely perceive the complex constraints of VNE, we propose a constraint-aware graph representation method tailored for VNE. Specifically, we design a heterogeneous modeling module for cross-graph status interactions. We also devise several feasibility-consistency augmentations and utilize contrastive learning to bring representations under different views close, which enhances the sensitivity of policy towards path-bandwidth constraints. Our contributions are summarized as follows.
  • We propose a new CMDP modeling approach with constraint violation tolerance for VNE, which precisely evaluate the quality of solution and the degree of constraint violation.
  • We present a reachability-guided optimization objective to achieve persistent zero constraint violation required by VNE. Further to enhance the stability of policy optimization, we propose a novel adaptive reachability budget method that dynamically decides budgets.
  • We propose a constraint-aware graph representation method tailored for VNE, which consists of a path-bandwidth contrast module with feasibility-consistency augmentations to perceive connectivity and a heterogeneous modeling module for cross-graph status fusion.
  • We conduct extensive experiments in various network scenarios, showing the CONAL’s superiority on performance, training stability, generalization, scalability and practicability.
@@ 2 Problem Definition
System Modeling. In real-world network systems, as illustrated in Figure 1, user network services are virtualized as VN requests that continuously seek resources from the PN. Each arrived VN request, along with the current situation of the PN, constitutes an instance $I$, and we collect all such instances with a set $\mathcal{I}$. For each instance, $I=(\mathcal{G}_{v},\mathcal{G}_{p})\in\mathcal{I}$, where the PN $\mathcal{G}_{p}$ and VN $\mathcal{G}_{v}$ are modeled as undirected graphs, $\mathcal{G}_{p}=(N_{p},L_{p})$ and $\mathcal{G}_{v}=(N_{v},L_{v},\omega,\varpi)$, respectively. Here, $N_{p}$ and $L_{p}$ denote the sets of physical nodes and links, indicating servers and their interconnections; $N_{v}$ and $L_{v}$ denote the sets of virtual nodes and links, representing services and their relationships; $\omega$ and $\varpi$ denote the arrival time and lifetime of VN request. We denote $\mathcal{C}(n_{p})$ as the computing resource availability for the physical node $n_{p}\in N_{p}$, and $\mathcal{B}(l_{p})$ as the bandwidth resource availability of the physical link $l_{p}\in L_{p}$. Besides, $\mathcal{C}(n_{v})$ and $\mathcal{B}(l_{v})$ denote the demands for computing resource by a virtual node $n_{v}\in N_{v}$ and bandwidth resource by a virtual link $l_{v}\in L_{v}$. Like most VNE studies (Wu et al., 2024), we focus mainly on bandwidth attributes in links, the generalizable critical bottleneck in various network scenarios, ensuring that the proposed method is universal and extensible.
@@
Figure 1: An brief example of VNE problem. In the network system, VN requests arrive sequentially at the infrastructure to require the resources of PN. For the VNE instance, embedding a VN to the PN consists of node and link mapping processes, while considering intricate and hard constraints.
@@ 3 Methodology
To optimize resource utilization while guaranteeing solution’s feasibility, we formulate solution construction of each VNE instance as a CMDP (Altman, 2021). However, VNE’s hard constraints make constructing complete solutions difficult, hampering the precise assessment of solution quality and the extent of constraint violations. Thus, we propose a violation-tolerant mapping method to ensure complete solution construction and customize a measurement function to evaluate violations.
CMDP for VNE. We consider each decision as identifying a proper physical node $n_{p}$ for each virtual node $n_{v}$ until all virtual nodes of VN are placed. Like existing works (Yan et al., 2020; Zhang et al., 2022; 2023b), due to the large combinatorial space of link mapping process, we incorporate link mapping into the state transitions, i.e., routing the prepared incident links $\delta^{\prime}(n_{v})$ of virtual node $n_{v}$.
Definition 1 (Prepared incident links of virtual node).
Let $\mathcal{N}_{v}^{t}$ denote the set of virtual nodes that have already been placed, and $n_{v}^{t}$ denote the to-be-placed virtual node at decision timestep $t$. We define $\delta(n_{v}^{t})$ as the set of virtual links incident to $n_{v}^{t}$. For each virtual link $l_{v}\in\delta(n_{v}^{t})$, if the link’s opposite endpoint $n_{v}^{\prime}$ is already placed, i.e., $n_{v}^{\prime}\in\mathcal{N}_{v}^{t}$, then we include $l_{v}$ in a subset $\delta^{\prime}(n_{v}^{t})$. The subset $\delta^{\prime}(n_{v}^{t})$ consists of what we term as the prepared incident links of the virtual node $n_{v}^{t}$. These links are considered prepared because, upon the placement of $n_{v}^{t}$, both endpoints of each link in $\delta^{\prime}(n_{v}^{t})$ are placed, necessitating the routing of these links. See Appendix D.1 for an example.
@@
Subsequently, we route all prepared incident links $\delta^{\prime}(n_{v}^{t})$ of virtual node $n_{v}^{t}$ sequentially. For each virtual link $l_{v}\in\delta^{\prime}(n_{v}^{t})$, we use the $k$-shortest path algorithm to find a set of physical paths. If there are available paths that do not violate constraints, we select the shortest one that consumes the least resources. In this case, we consider the violation as $H_{L}(l_{v})=\max_{l_{p}\in f_{L}(l_{v})}(\mathcal{B}(l_{v})-\mathcal{B}(l_{p}))$. However, if all paths violate constraints, we calculate the extent of the constraint violation for each path. The path with the least amount of violation is then selected to route the virtual link, whose violation is calculated by $H_{L}(l_{v})=\sum_{l_{p}\in f_{L}(l_{v})}(\mathcal{B}(l_{v})-\mathcal{B}(l_{p}))$. After $\delta^{\prime}(n_{v}^{t})$ are routed, we define the bandwidth violations as $h_{L}^{t}=H_{L}(s_{t+1})=\max_{l_{v}\in\delta^{\prime}(n_{v}^{t})}H_{L}(l_{v})$. During training, we always set the solution feasibility flag $\varkappa$ to $1$ to measure the final R2C metric until the solution is complete.
@@ 3.2 Reachability-guided Optimization with Adaptive Budget
Standard CMDPs focus on optimizing discounted cumulative costs to meet long-term safety, which fails to meet the consistent satisfaction requirements of VNE in all states (Liu et al., 2021). To guarantee the solution feasibility of VNE, we consider reachability analysis into CMDP to achieve state-wise zero-violation optimization (Yu et al., 2022). This objective significantly expands the feasible set of policies and mitigates the conservativeness of the policy. Additionally, to enhance the stability of policy optimization, we propose a novel adaptive reachability budget method that dynamically decides budgets, rather than always zero budget for unsolvable instances. For policy training, we leverage the Lagrange version of the PPO algorithm (Ray et al., 2019).
@@
$\max_{\pi}\mathbb{E}_{s}\left[V_{r}^{\pi}(s)\right]\text{, s.t. }\mathbb{E}_{s}\left[V_{c}^{\pi}(s)\right]\leq 0,$ (2)
However, since the non-negativity of $C$, the optimization focuses on constraint satisfaction rather than reward maximization, which causes a highly conservative policy. Thus, we consider the Hamilton-Jacobi (HJ) reachability analysis (Bansal et al., 2017) into the CMDP to obtain a policy with the best possible performance and least violations, improving both quality and feasibility of VNE solutions. We need the following concepts to further present our optimization objective.
Definition 2 (Feasible value function).
@@
Definition 3 (Feasible region).
@@
$\max_{\pi}\mathbb{E}_{s}\left[V^{\pi}_{r}(s)\cdot\mathbb{I}_{s\in\mathcal{S}_{f}}-V_{h}^{\pi}(s)\cdot\mathbb{I}_{s\notin\mathcal{S}_{f}}\right],\text{ s.t. }\mathbb{E}_{s}\left[V_{h}^{\pi}(s)\right]\leq 0,\forall s\in\mathcal{S}_{f},$ (3)
where $\mathbb{I}$ is the indicator function. Compared to the problem (2), our reachability-guided optimization problem (3) finds the largest feasible sets, bringing less conservativeness and better performance.
To solve the problem (3), we reformulate it in the Lagrangian version as follows:
$\min_{\lambda}\max_{\pi}\mathbb{E}_{s}\left[V^{\pi}_{r}(s)\cdot\mathbb{I}_{s\in\mathcal{S}_{f}}-V_{h}^{\pi}(s)\cdot\mathbb{I}_{s\notin\mathcal{S}_{f}}+\lambda V_{h}^{\pi}(s)\cdot\mathbb{I}_{s\in\mathcal{S}_{f}}\right].$ (4)
@@
$\forall s\in\tau,D^{\pi^{\prime}}_{h}(s)=\max_{s^{\prime}\in\tau^{\prime}}C(s^{\prime}).$ (5)
During the training process, we update the surrogate policy $\pi^{\prime}\leftarrow\pi$ over multiple iterations. With the reachability guidance of $\pi^{\prime}$, the main policy $\pi$ gradually improves constraint-aware decision-making, enhancing the stability of training. Through this iterative learning, both policies $\pi$ and $\pi^{\prime}$ achieve better constraint satisfaction while maintaining the exploration for better solutions. Finally, we obtained the refined Lagrangian objective with the adaptive reachability budget as follows:
$\min_{\lambda}\max_{\pi}\mathbb{E}_{s}\left[V^{\pi}_{r}(s)-\lambda\left(V_{h}^{\pi}(s)-D^{\pi^{\prime}}_{h}(s)\right)\right].$ (6)

### v2

# Towards Constraint-aware Learning for Resource Allocation in NFV Networks

@@ Abstract
Virtual Network Embedding (VNE) is a fundamental resource allocation challenge that is associated with hard and multifaceted constraints in network function virtualization (NFV). Existing works for VNE struggle to handle such complex constraints, leading to compromised system performance and stability. In this paper, we propose a CONstraint-Aware Learning framework, named CONAL, for efficient constraint handling in VNE. Concretely, we formulate the VNE problem as a constrained Markov decision process with violation tolerance, enabling precise assessments of both solution quality and constraint violations. To achieve the persistent zero violation to guarantee solutions’ feasibility, we propose a reachability-guided optimization with an adaptive reachability budget method. This method also stabilizes policy optimization by appropriately handling scenarios with no feasible solutions. Furthermore, we propose a constraint-aware graph representation method to efficiently learn cross-graph relations and constrained path connectivity in VNE. Finally, extensive experimental results demonstrate the superiority of our proposed method over state-of-the-art baselines. Our code is available at https://github.com/GeminiLight/conal-vne.
Reinforcement Learning, Resource Allocation, Network Function Virtualization
@@ 1 Introduction
Network Function Virtualization (NFV) is a promising technique that facilitates the deployment of multiple Virtual Networks (VNs) tailored to user network demands within a shared Physical Network (PN) infrastructure (Yi et al., 2018). It is vital for domains such as cloud computing, edge computing and 5G, where dynamic and efficient resource management is essential (Zhuang et al., 2020). Virtual Network Embedding (VNE), a fundamental resource allocation problem in NFV, is critical for maintaining high Quality of Service (QoS). This process, involving the mapping of VNs to PNs, represents a significant challenge. It is an NP-hard Combinatorial Optimization Problem (COP) characterized by intricate and hard constraints (Rost & Schmid, 2020).
@@
Traditional solutions to VNE ranging from exact to heuristic methods often suffer from either expensive computation times or limited performance in complex network scenarios (Zhang et al., 2018; Fan et al., 2023). Recently, Reinforcement Learning (RL) has been a potential direction for VNE, which learns effective solving policies without relying on high-quality labeled data. Typically, existing RL-based methods solve the VNE problem as a Markov Decision Process (MDP)  (Haeri & Trajković, 2017; Yan et al., 2020; Zhang et al., 2024b), where a neural policy is iteratively optimized through interactions with an environment. They construct VNE solutions by sequentially selecting an available physical node to place each virtual node, until the solution is completed or constraints are violated. However, the hard constraints of VNE require a persistent zero-violation at each decision timestep, which results in numerous failure samples where constraints are violated during training. For failure samples, existing studies (e.g., Yao et al. (2020); Zhang et al. (2022)) consider them as noisy and only train with data that violate no constraint; or others, like Yan et al. (2020) and He et al. (2023a), consider fixed penalties to them and discourage violations.
@@
Although these RL-based algorithms (He et al., 2023a; Zhang et al., 2024b) have shown efficacy, they still suffer from several significant problems in handling complex constraints of VNE. Firstly, ignoring failure samples makes policies prone to violating critical constraints, while employing fixed penalties in reward signals does not accurately reflect the severity of constraint violations. Thus, these methods underestimate valuable sample information and hamper the learning of constraint-aware policies, which results in low feasibility guarantees. Secondly, it is hard to avoid to encounter unsolvable instances whose feasible sets are empty in practical scenarios, due to insufficient physical resource availability or excessive virtual resource requests. It is impractical to distinguish solvable and unsolvable instances, since checking the instance solvability of an NP-hard problem is time-consuming. Failure samples caused by unsolvable instances further complicate the constraint learning process, and negatively impact the stability of training and policy performance. See Appendix C for a preliminary study highlighting their negative impact on training stability. Thirdly, VNE constraints are complex and multifaceted, involving cross-graph status interactions and bandwidth-constrained path connectivity, which are not adequately captured by exsiting feature extractors.
@@
To address these challenges, we propose a CONstraint-Aware Learning framework for VNE, named CONAL, achieving high solution feasibility guarantee and training stability. Concretely, to optimize performance while enhancing constraint satisfaction, we formulate the VNE problem as Constrained MDP (CMDP) (Altman, 2021) with violation-tolerance. Although constraint violations often trigger early termination during solution construction, leading to incomplete solutions and complicating the measurement of both solution quality and violation degree, this method enables complete solution construction and precise measurement. Additionally, we introduce a reachability analysis into the optimization objective to achieve persistent zero constraint violation required by VNE. To handle the existence of unsolvable instances whose constraints are impossibly satisfied, we propose an adaptive reachability budget method to make the policy optimization more stable. It dynamically decides the violations caused by a surrogate policy as budgets based on the instance’s solvability, rather than always setting budgets to zero. Furthermore, to finely perceive the complex constraints of VNE, we propose a constraint-aware graph representation method tailored for VNE, which includes a heterogeneous modeling module for cross-graph status interactions. We also devise several feasibility-consistency augmentations and utilize contrastive learning to bring representations under different views close, which enhances the sensitivity of policy towards path-bandwidth constraints.
Our main contributions are summarized as follows: (1) We propose a violation-tolerant CMDP modeling approach for VNE, which precisely evaluates the quality of solution and the degree of constraint violation. To achieve persistent zero constraint violation, we also present a reachability-guided optimization objective. (2) We present a pioneering work on the study of unsolvable instances negatively impacting the stability of policy optimization in COP, and propose a novel adaptive reachability budget method to handle such challenges. (3) We propose a constraint-aware graph representation method tailored for VNE, where a path-bandwidth contrast module with feasibility-consistency augmentations perceives connectivity at a fine-grained level. (4) We conduct extensive experiments in various network scenarios, showing the CONAL’s superiority on performance, training stability, generalization, scalability and practicability.
@@ 2 Problem Definition
System Modeling. In real-world network systems, as illustrated in Figure 1, user network services are virtualized as VN requests that continuously seek resources from the PN. Each arrived VN request, along with the current situation of the PN, constitutes an instance $I$, and we collect all such instances with a set $\mathcal{I}$. For each instance, $I=(\mathcal{G}_{v},\mathcal{G}_{p})\in\mathcal{I}$, where the PN $\mathcal{G}_{p}$ and VN $\mathcal{G}_{v}$ are modeled as undirected graphs, $\mathcal{G}_{p}=(N_{p},L_{p})$ and $\mathcal{G}_{v}=(N_{v},L_{v},\omega,\varpi)$, respectively. Here, $N_{p}$ and $L_{p}$ denote the sets of physical nodes and links, indicating servers and their interconnections; $N_{v}$ and $L_{v}$ denote the sets of virtual nodes and links, representing services and their relationships; $\omega$ and $\varpi$ denote the arrival time and lifetime of VN request. We denote $\mathcal{C}(n_{p})$ as the computing resource availability for the physical node $n_{p}\in N_{p}$, and $\mathcal{B}(l_{p})$ as the bandwidth resource availability of the physical link $l_{p}\in L_{p}$. Besides, $\mathcal{C}(n_{v})$ and $\mathcal{B}(l_{v})$ denote the demands for computing resource by a virtual node $n_{v}\in N_{v}$ and bandwidth resource by a virtual link $l_{v}\in L_{v}$.
@@
Figure 1: A brief example of VNE problem. In the network system, VN requests arrive sequentially at the infrastructure to require the resources of PN. For the VNE instance, embedding a VN to the PN consists of node and link mapping processes, while considering intricate and hard constraints.
@@ 3 Methodology
To optimize resource utilization while guaranteeing solution’s feasibility, we formulate solution construction of each VNE instance as a CMDP (Altman, 2021). However, VNE’s hard constraints make constructing complete solutions difficult, hampering the precise assessment of solution quality and constraint violations. Thus, we propose a violation-tolerant mapping method to ensure complete solution construction and a measurement function to evaluate violations.
CMDP for VNE Solving. We consider each decision as identifying a proper physical node $n_{p}$ for each virtual node $n_{v}$ until all virtual nodes of VN are placed. Like existing works (Yan et al., 2020; Zhang et al., 2022, 2024b), due to the large combinatorial space of link mapping process, we incorporate link mapping into the state transitions, i.e., routing prepared incident links $\delta^{\prime}(n_{v})$ of virtual node $n_{v}$.
Definition 3.1 (Prepared incident links of virtual node).
Let $\mathcal{N}_{v}^{t}$ denote the set of virtual nodes that have already been placed, and $n_{v}^{t}$ denote the to-be-placed virtual node at decision timestep $t$. We define $\delta(n_{v}^{t})$ as the set of virtual links incident to $n_{v}^{t}$. For each virtual link $l_{v}\in\delta(n_{v}^{t})$, if the link’s opposite endpoint $n_{v}^{\prime}$ is already placed, i.e., $n_{v}^{\prime}\in\mathcal{N}_{v}^{t}$, then we include $l_{v}$ in a subset $\delta^{\prime}(n_{v}^{t})$. The subset $\delta^{\prime}(n_{v}^{t})$ consists of what we term as the prepared incident links of the virtual node $n_{v}^{t}$. These links are considered prepared because, upon the placement of $n_{v}^{t}$, both endpoints of each link in $\delta^{\prime}(n_{v}^{t})$ are placed, necessitating the routing of these links. See Appendix D.1 for an illustrative example.
@@
Subsequently, we route all prepared incident links $\delta^{\prime}(n_{v}^{t})$ of virtual node $n_{v}^{t}$ sequentially. For each virtual link $l_{v}\in\delta^{\prime}(n_{v}^{t})$, we use the $k$-shortest path algorithm to find a set of physical paths. If there are available paths that do not violate constraints, we select the shortest one that consumes the least resources. In this case, we consider the violation as $H_{L}(l_{v})=\max_{l_{p}\in f_{L}(l_{v})}(\mathcal{B}(l_{v})-\mathcal{B}(l_{p}))$. However, if all paths violate constraints, we calculate the extent of the constraint violation for each path. The path with the least amount of violation is then selected to route the virtual link, whose violation is calculated by $H_{L}(l_{v})=\sum_{l_{p}\in f_{L}(l_{v})}(\mathcal{B}(l_{v})-\mathcal{B}(l_{p}))$. After $\delta^{\prime}(n_{v}^{t})$ are routed, we define the bandwidth violations as $h_{L}^{t}=H_{L}(s_{t+1})=\max_{l_{v}\in\delta^{\prime}(n_{v}^{t})}H_{L}(l_{v})$. During training, we always set the solution feasibility flag $\varkappa$ to 111 to measure the final R2C metric until the solution is complete.
@@ 3.2 Reachability-guided Optimization with Adaptive Budget
Standard CMDPs focus on optimizing discounted cumulative costs to meet long-term safety, which fails to meet the consistent satisfaction requirements of VNE in all states (Liu et al., 2021). To guarantee the solution feasibility of VNE, we consider reachability analysis into CMDP to achieve state-wise zero-violation optimization (Yu et al., 2022). This objective significantly expands the feasible set of policies and mitigates the conservativeness of the policy. Additionally, to enhance the stability of policy optimization, we propose a novel adaptive reachability budget method that dynamically decides budgets, rather than always zero budget for unsolvable instances. For policy training, we adopt the Lagrange version of the PPO algorithm (Ray et al., 2019).
@@
$\max_{\pi}\mathbb{E}_{s}\left[V_{r}^{\pi}(s)\right]\text{, s.t. }\mathbb{E}_{s}\left[V_{c}^{\pi}(s)\right]\leq 0,\vskip-4.0pt$ (2)
However, since the non-negativity of $C$, the optimization focuses on constraint satisfaction rather than reward maximization, causing a highly conservative policy. Thus, we consider the Hamilton-Jacobi (HJ) reachability analysis (Bansal et al., 2017) into the CMDP to obtain a policy with the best possible performance and least violations, improving both quality and feasibility of VNE solutions. We need the following concepts to present our optimization further.
Definition 3.2 (Feasible value function).
@@
Definition 3.3 (Feasible region).
@@
$\max_{\pi}$ $\mathbb{E}_{s}\!\Bigl{[}V^{\pi}_{r}(s)\,\mathbb{I}_{s\in\mathcal{S}_{f}}\;-\;V_{h}^{\pi}(s)\,\mathbb{I}_{s\notin\mathcal{S}_{f}}\Bigr{]},$ (3)
[…]
where $\mathbb{I}$ is the indicator function. Compared to the problem (2), our reachability-guided optimization problem (3) finds the largest feasible sets, bringing less conservativeness and better performance. To solve the problem (3), we reformulate it in the Lagrangian version as follows:
$\min_{\lambda}\max_{\pi}\mathbb{E}_{s}\Bigl{[}V^{\pi}_{r}(s)\cdot\mathbb{I}_{s\in\mathcal{S}_{f}}-V_{h}^{\pi}(s)\cdot\mathbb{I}_{s\notin\mathcal{S}_{f}}+\lambda V_{h}^{\pi}(s)\cdot\mathbb{I}_{s\in\mathcal{S}_{f}}\Bigr{]}$
@@
$\forall s\in\tau,D^{\pi^{\prime}}_{h}(s)=\max_{s^{\prime}\in\tau^{\prime}}C(s^{\prime}).\vskip-6.0pt$ (5)
During the training process, we update the surrogate policy $\pi^{\prime}\leftarrow\pi$ over multiple iterations. With the reachability guidance of $\pi^{\prime}$, the main policy $\pi$ gradually improves constraint-aware decision-making, enhancing the training stability. Through this iterative learning, both policies $\pi$ and $\pi^{\prime}$ achieve better constraint satisfaction while maintaining exploring for better solutions. Finally, we obtain the refined Lagrangian objective with adaptive reachability budget:
$\min_{\lambda}\max_{\pi}\mathbb{E}_{s}\left[V^{\pi}_{r}(s)-\lambda\left(V_{h}^{\pi}(s)-D^{\pi^{\prime}}_{h}(s)\right)\right].\vskip-10.0pt$ (6)